# Post-detection inference for sequential changepoint localization

### v3

@@ Appendix A Omitted proofs
For all $t\in\mathbb{N},\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}^{*}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\geq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t].$ The last step follows from Assumption 3.3. The rest of the proof is the same as in the proof of Theorem 3.2. ∎
@@
First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $[\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{T,L})}{P_{\infty}(\tau\geq T)}.$ (A.3)
@@
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (A.4)
Therefore, combining (A.3) and (A.4), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.5)
@@
the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (A.6)
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{o}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})\mid r_{t}^{*}))$
@@
the last equality holds because the event $[\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
@@
Appendix B Computationally and statistically efficient subroutines in Algorithm 2
@@
$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (B.1)
@@
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (B.2)
@@
$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (B.3)
@@
$M_{t}=\begin{cases}\frac{L_{\hat{T}}(\hat{\theta}_{1,\hat{T}:\tau};X_{1},\cdots,X_{\tau})}{L_{t}(\hat{\theta}_{1,t:\tau};X_{1},\cdots,X_{\tau})},&\text{if }t\leq\tau<\infty,\\ -\infty,&\text{if }t>\tau,\text{ or }\tau=\infty.\\ \end{cases}$ (B.4)
@@
$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (B.5)
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$. In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
@@
$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (B.6)
@@
$V_{t,\infty}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j}\text{ or }t_{1}^{j}=\infty,\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{otherwise,}\\ \end{cases}$ (B.7)
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section B.1). We define the confidence set as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (B.8)
@@
Corollary B.1.
$\mathcal{C}^{\prime}_{1}$ defined in (B.8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.4.
Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section G.3.
B.1 A concrete example: Gaussian mean shift problem
@@
$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (B.9)
@@
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (B.10)
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (B.4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark B.2) and they are finite. So, we can have $L=\infty$.
@@
Remark B.2.
[…]
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (B.11)
[…]
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section G.2.
The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology. It is worth noting that $C^{\prime}$, as defined in (B.8), is practically implementable as long as the following relatively weak conditions are met:
@@
    If we employ the likelihood-ratio-based test statistic (B.4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
@@
B.2 Another concrete example: Poisson rate change problem

### v4

@@ Appendix A Omitted proofs
Since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$ and is independent of data, we have $\forall t\in\mathbb{N}$,
$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/(\alpha r_{t}^{*})]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/(\alpha r_{t}^{*})\mid r_{t}^{*}]]\leq\mathbb{E}[\alpha r_{t}^{*}]=\alpha\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]\leq\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t],$ $\forall F_{0}\in\mathcal{P}_{0}$.
The last steps follow from Assumption 3.5. The rest of the proof is the same as in the proof of Theorem 3.2. ∎
@@
It follows from the definition of the least favorable distribution that
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right).$
Therefore, for any $t\leq T-1$, using the above inequality and the assumption (3.10a) we get
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
$\leq\frac{1}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right)$
$\leq\frac{1}{\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)}\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}^{*}(X_{i})}{f_{1}^{*}(X_{i})}\leq 2/(\alpha r_{t})\right).$
Similarly, for any $t\geq T+1$,
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha r_{t}),\tau\geq t\mid\tau\geq T\right)$
$\leq\frac{1}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha r_{\tau})\right)$
$\leq\frac{1}{\mathbb{P}_{F_{0},\infty}(\tau\geq T)}\left[\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)+\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\right]$
$\leq\frac{1}{\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)}\left[\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)+\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)\right].$
In the last step, $\mathbb{P}_{F_{0},\infty}(\tau\geq T)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)$ follows from (3.10a), $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ follows from the fact that $(F_{0}^{*},F_{1}^{*})$ is an LFD pair, and $\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)$ follows from (3.10c).
The remaining calculations are the same as those done in the proof of the previous theorem, with $f_{i}$ replaced by $f_{i}^{*}.$
Proof of Theorem 4.1.
First note the fact that $\mathbb{P}_{F_{0},T,F_{1}}(\tau<T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau<T),$ since the event $\tau<T]=\displaystyle\cup_{i=1}^{T-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq T\}$.
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\frac{P_{T}(M_{T}>C_{T,L})}{P_{\infty}(\tau\geq T)}.$ (S6)
@@
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (S7)
Therefore, combining (S6) and (S7), we obtain $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha.$
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (S8)
@@
the last equality holds because the event $\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(M_{t}\leq\operatorname{Quantile}(1-c;M_{t},\{M_{t,\infty}^{j}(\theta_{0},\theta_{1})\}_{j=1}^{B}))\geq\frac{\lceil(1-c)(B+1)\rceil}{B+1}\geq 1-c,\text{ for all }t.$ (S9)
@@
$\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{0}\in\mathcal{S}_{t-1})$ $=\mathbb{E}(\mathbb{P}_{\theta_{0},t,\theta_{1}}(\theta_{1}\in\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})\mid r_{t}^{*}))$
@@
the last equality holds because the event $\tau<T]$ does not depend $X_{i}$ for $i\geq T$. Now,
@@
A.1 Auxilary Lemmas
Lemma A.2.
Define, $r_{T}=\frac{1+\sum_{i=1}^{N}\mathds{1}(\tau\geq T)}{N+1}$. Then, for fixed $s\in(0,1)$, $\{r_{T}^{-s}\}_{N\geq 1}$ is uniformly integrable.
Proof.
[…]
$\mathbb{E}[r_{T}^{-1}]=(N+1)\mathbb{E}\left[\frac{1}{K_{N}+1}\right]=\frac{1-(1-p_{T})^{N+1}}{p_{T}}\leq\frac{1}{p_{T}}.$
This bound $\frac{1}{p_{T}}$ is a constant that does not depend on $N$. Hence, $r_{t}^{-s}$ is uniformly integrable. ∎
Lemma A.3.
Suppose, $\tau_{1},\cdots,\tau_{N}$ are i.i.d. random variables. Define, $r_{t}=\frac{1+\sum_{i=1}^{N}\mathds{1}(\tau_{i}\geq t)}{N+1}$, for $t\in\mathbb{N}.$ If each $\tau_{j}$ is stochastically larger than a random variable $\tau$, then $r_{\tau}$ is a p-value.
Proof.
Suppose $\tau_{N+1}$ is independent and identically distributed (i.i.d.) with $\tau_{1},\cdots,\tau_{N}$. By Theorem 1.A.2 of [26], $\tau_{N+1}$ is stochastically larger than $\tau$ implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\tau\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\tau_{N+1}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$. By direct calculation,
$\mathbb{P}[r_{\tau}\leq(1+k)/(N+1)]=\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\tau)\leq k]$ $=\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\psi_{1}(Z))\leq k]$
$\leq\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\psi_{2}(Z))\leq k]$
$=\mathbb{P}[\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq\tau_{N+1})\leq k]\leq\frac{1+k}{N+1},$
for all $k=0,1,\cdots,N.$ Therefore, $r_{\tau}$ is a p-value. ∎
Appendix B Subtleties regarding Subroutine $\mathcal{L}$ in Algorithm 1
The problem with $L=\infty$ is that the stopping time may never be reached for some sampled sequence. One can set $L=\infty$ if $\mathcal{A}$ is known to satisfy $\mathbb{P}_{T}(\tau<\infty)=1$, for all $T\in\mathbb{N}$ (if there is a change, $\mathcal{A}$ stops almost surely). This is a very mild assumption because any reasonable change detector should eventually raise an alarm if there is a change; indeed, for example, this holds for all algorithms with finite ARL. If $\mathcal{A}$ is not known to satisfy this condition, one can simply pick $L<\infty$ (The value of $L$ does not need to be a fixed constant; it can be set adaptively depending on the observed data sequence, for instance, one can choose $L$ to be $2\tau$.) If $L<\infty$ and the $j$th sequence stopped because $\mathcal{A}$ raised an alarm at or before $L$, then we define $M^{j}_{t,L}$ as usual. If $\mathcal{A}$ had not stopped by time $L$, we then define $M^{j}_{t,L}=\infty$ to ensure a lower bound on the original test statistic (if we could compute it). For $L<\infty$,
[…]
@@ Appendix C Computationally and statistically efficient subroutines in Algorithm 2
$L_{t}(\theta_{1};X_{1},\cdots,X_{n})=\prod_{i=1}^{t-1}f_{\theta_{0}}(X_{i})\prod_{i=t}^{n}f_{\theta_{1}}(X_{i}).$ (S1)
@@
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})}.$ (S2)
@@
$M_{t}^{(n)}(X_{1},\cdots,X_{n})=\displaystyle\max_{1\leq i\leq n}\frac{L_{i}(\hat{\theta}_{1,i:n};X_{1},\cdots,X_{n})}{L_{t}(\hat{\theta}_{1,t:n};X_{1},\cdots,X_{n})}.$ (S3)
@@
[…]
@@
$Y_{n}^{t}(x,\theta):=\begin{cases}G(x,\theta_{0}),&\text{if }n<t\\ G(x,\theta),&\text{ otherwise. }\end{cases}$ (S5)
Let us provide a few examples of the function $G$. If the distributions in the parametric family are continuous, the inverse distribution function $F_{\theta}^{-1}$ exists, and it is easy to verify that $G(x,\theta)=F_{\theta}^{-1}(F_{\theta^{\prime}_{t}}(x))$ works. In the case of a location family, a simple transformation is $G(x,\theta)=x+\theta-\theta^{\prime\prime}$ and similarly, for scale family, we have $G(x,\theta)=x\times\theta/\theta^{\prime\prime}$. In the case of discrete distributions, even if a suitable $G$ is not directly available, generating such sequences $\{Y_{n}^{t}(x,\theta)\}_{n}$ across all $\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$ may still be possible, as demonstrated in Appendix C of the Supplementary Material for the Poisson distribution.
@@
$V_{t,L}^{j}:=\begin{cases}-\infty,&\text{if }t>t_{2}^{j},\\ \displaystyle\max_{\max\{t,t_{1}^{j}\}\leq t^{\prime}\leq t_{2}^{j}}\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}M_{t}^{(t^{\prime})}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta);t^{\prime}),&\text{if }t\leq t_{2}^{j}\leq L,\\ \infty,&\text{if }t\leq t_{2}^{j}\text{ and }t_{2}^{j}>L\\ \end{cases}.$ (S6)
@@
[…]
which are often straightforward to compute even with $L=\infty$ (e.g., in a Gaussian setting with CUSUM or SR-type change detectors, see Section C.1). We define the confidence set as
$\mathcal{C}^{\prime}_{1}=\{t\in\{1,\cdots,\tau\}:M_{t}\leq\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},V_{t,L}^{1},\cdots,V_{t,L}^{B})\}\text{ and}$ (S8)
@@
Corollary C.1.
$\mathcal{C}^{\prime}_{1}$ defined in (S8) satisfies the same coverage guarantees as $\mathcal{C}_{1}$ in Theorem 4.3.
Next, we present a concrete example of the Gaussian mean shift problem to clarify our method. We illustrate another example with the Poisson distribution in Section H.4.
C.1 A concrete example: Gaussian mean shift problem
@@
$\mathcal{S}^{\prime}_{\tau-t+1}=\left[\bar{X}_{t:\tau}-\frac{s_{\tau-t+1}(\beta)}{\sqrt{\tau-t+1}},\bar{X}_{t:\tau}+\frac{s_{\tau-t+1}(\beta)}{\sqrt{{\tau-t+1}}}\right]\cap\Theta_{1},$ (S9)
@@
$Y_{n}^{t}(X_{n}^{j},\theta):=\begin{cases}\epsilon_{n}^{j}+\theta_{0},\text{ if }n<t,\\ \epsilon_{n}^{j}+\theta,\text{ otherwise, }\end{cases}$ (S10)
where $\epsilon_{n}^{j}:=X_{n}^{j}-\theta_{0}\sim N(0,1)$. We consider the likelihood-ratio-based test statistic $M_{t}(X_{1},\cdots,X_{\tau})$, as defined in (S4), with $\hat{\theta}_{1,t:n}=\frac{1}{n-t+1}\sum_{k=t}^{n}X_{k}$. Let, $\tau^{\theta}_{j}$ is the detection time for the sequence $\{Y_{n}^{t}(X_{n}^{j},\theta)\}_{n}$ and $M_{t,\infty}^{j}(\theta)=M_{t}(Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{\tau^{\theta}_{j}}^{t}(X_{\tau^{\theta}_{j}}^{j},\theta))$. Suppose, we can find $t_{1}^{j},t_{2}^{j}\in\mathbb{N}$, such that $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}$ (these are easy to compute for specific detection algorithms, such as CUSUM or SR type detectors, as we explain in Remark C.2) and they are finite. So, we can have $L=\infty$.
@@
Remark C.2.
For weighted CUSUM or SR detectors, finding $t_{1}^{j},t_{2}^{j}$ is often straightforward. For some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$, these are defined as
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (S11)
 $\text{ and }\penalty 10000\ \tau_{\text{wsr}}=\inf\left\{n\in\mathbb{N}:\sum_{j=1}^{n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}.$ (S12)
For example, consider $\mathcal{P}_{1}=\{N(\theta,1):\theta\geq\theta^{*}\}$, for some $\theta^{*}\geq\theta_{0}$. Observe that for all $\theta\in\Theta_{1}$, $\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}=\exp\{\frac{1}{2}(\theta-\theta_{0})\times\sum_{i=j}^{n}(2X_{i}-\theta_{0}-\theta)\}$ increases as $X_{i}$’s increases and hence, both $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ decreases as $X_{i}$’s increases. Therefore, $t_{1}^{j}=\inf_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\tau^{\theta}_{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Similarly, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\leq\theta^{*}\}$, for some $\theta^{*}\leq\theta_{0}$, we have $t_{1}^{j}=\tau^{\inf\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$ and $t_{2}^{j}=\tau^{\sup\mathcal{S}^{\prime}_{\tau-t+1}}_{j}$. Also, for $\mathcal{P}_{1}=\{N(\theta,1):\theta\neq\theta_{0}\}$, if we employ Gaussian weights (it is conjugate prior, so closed-form expressions for $\tau_{\text{wcusum}}$ and $\tau_{\text{wsr}}$ can be found), one can obtain similar forms of $t_{1}^{j},t_{2}^{j}$, depending on the mean of the weight distribution. One can check that similar results extend to other parametric families, such as the Gaussian/Laplace/exponential scale change problems with weighted CUSUM and SR detectors, as well as the PFA-controlling change detectors used in Section H.3.
The Gaussian mean shift example we provided serves as a concrete and illustrative case to clarify the methodology. It is worth noting that $C^{\prime}$, as defined in (S8), is practically implementable as long as the following relatively weak conditions are met:
@@
    If we employ the likelihood-ratio-based test statistic (S4), then we must be able to maximize $L_{i}(\hat{\theta}_{1,i:t^{\prime}};Y_{1}^{t}(X_{1}^{j},\theta),\cdots,Y_{t^{\prime}}^{t}(X_{t^{\prime}}^{j},\theta))$ over $\theta$ on a given set (either exactly or numerically).
@@
C.2 Another concrete example: Poisson rate change problem